OpenStudy (anonymous):

Find the solutions to the systems of nonlinear equations y=4x x^2+y^2=68

OpenStudy (anonymous):

putting y=4x in equation 2 we get \[x ^{2}+16x ^{2}=68\] and x^2=4; gives x=2,-2 and corresponding y=8,-8.